Operations Report – February 20th

Crew 205 Operations Report 20 – FEB – 2019

SOL: 10

Name of person filing report: Verónica Triviño

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): Turned on at 1400 (20thFeb), turned off at 0800 (20thFeb)

Solar— SOC % – Turned off Gen (20th Feb) 57%, Turned on Gen (20th Feb) 90%

Diesel Reading – 75%

Propane Reading – ~ 34%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 39%, ~ 214 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 189 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01414004 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 51.0 hours

Beginning charge: 100%

Ending charge: 49%

Currently charging: Yes

Curiosity rover used: Used

Hours: 75.3 hours

Beginning charge: 100%

Ending charge: 71%

Currently charging: Yes

Spirit rover used: Used

Hours: 76.1

Beginning charge: 100%

Ending charge: 53%

Currently charging: Yes

Notes on rovers: Spirit and Opportunity were not at 100% charge at the beginning of the first attempt at the EVA due to problems with water and the external breaker of RAM. Breaker was reset and is now functional. They were left to charge for an hour in order to reach 100% charge and were then fit to be used on EVA.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? N/A

Crew Car used and why, where? N/A

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Nothing to report

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report